OpenStudy (anonymous):

Alex is as old in years as his dog is in months. In one year the dog's age in months will be twice Alex's age in years. Find alex's age now

OpenStudy (anonymous):

Alex's age: A in years Dog's age: D in years A=12D (A+1)*2=12(D+1) so plugging in A from the first equation into the second you get 12D+2=6(D+1) so 6D+1=3D+3 so 3D=2 so D=2/3 or 8 months A=12(2/3)=8 years

OpenStudy (anonymous):

Ugh my brain has officially stopped functioning :l

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you're welcome!